# Open Wireless Digital Twin: End-to-End 5G Mobility Emulation with OpenAirInterface and Sionna RT

### v3

@@ 2 METHODS
This section outlines the architecture and operational workflow of the proposed platform. Subsection 2.1 describes the modular architecture of the OWDT platform, detailing how OSSs and open data are integrated to create a realistic and flexible 5G emulation environment. Subsection 2.2 explains the overall workflow of the OWDT platform, illustrating how data flows through the system from scenario setup and simulation to near-RT monitoring of KPIs.
2.1 SYSTEM ARCHITECTURE
@@
We realized the OWDT by integrating the OAI, which reproduces the 5G NR protocol stack from the PHY layer and above, with NVIDIA Sionna RT v0.18.0, which enables the construction of a deterministic channel model through ray tracing. In this environment, CIRs pre-generated using Sionna RT were applied as propagation channels between the gNB and UE in the OAI. This allows us to simulate a scenario in which the UE moves within a 3D map while performing 5G communication, even though the gNB and UE remain physically connected over the cable. Furthermore, we can observe how KPIs, such as reference signal received powers (RSRPs), modulation and coding schemes (MCSs), block error rates (BLERs), and throughputs, dynamically change in the near-RT during the scenario. These measurements were performed using the O-RAN standard KPM v3.00 [38], which is one of the service models supported by FlexRIC [20], a near-RT RIC that provides an xApp SDK framework developed within OAI.
@@
The fundamental architecture of the OWDT is illustrated in the right side of Fig. 1. The hardware components that constitute the architecture are as follows: (i) A COTS server to implement the 5GCN, gNB, and near-RT RIC, along with an SDR device responsible for RF signal processing. The server was equipped with an AMD Ryzen 9 7900X 12-core processor (×24), 32 GiB RAM, and ran Ubuntu 22.04.4 LTS with Linux kernel version 6.8.1-060801-generic. An NI USRP B200 was used as the SDR device and connected via USB. In practice, this setup was realized using Allbesmart’s OAIBOX™ 40 [39]. (ii) A computation server used for ray-tracing-based radio propagation simulation with Sionna RT. The server was equipped with an Intel(R) Xeon(R) Silver 4310 @ 2.10 GHz CPU, 256 GiB RAM, an NVIDIA A100 GPU (80 GB), and ran Ubuntu 22.04.4 LTS. (iii) A COTS laptop and SDR device pair for implementing an NR UE, or alternatively, any UE module that can be physically connected to the gNB via an RF cable. In this study, the Quectel RM500Q-GL UE module was connected to a laptop with an 11th Gen Intel Core i7-11370H @ 3.30 GHz (×8), 15.4 GiB RAM, and Ubuntu 20.04 LTS via USB.
2.2 WORKFLOW
@@
2.2.1 Preparation of 3D Building Data
To achieve a higher resolution in radio propagation simulations, it is essential to prepare highly detailed 3D building data in terms of size and location and perform site-specific deterministic evaluations. The 3D urban data provided by PLATEAU, an open data project promoted by the Ministry of Land, Infrastructure, Transport, and Tourism (MLIT) in Japan [40], adopts building height measurements obtained via aerial surveying and generally provides higher accuracy in the height dimension than OpenStreetMap (OSM) [41]. These data are available in CityGML format, which cannot be directly handled by Sionna RT in this study. Therefore, the open-source 3D CG software Blender 4.2 [42] was used to convert it into the PLY format. If OpenStreetMap is used as 3D building data, it can be imported into Blender using Blosm [43]. The Mitsuba Blender Add-on [44] was used to assign materials to the polygon surfaces of the imported 3D model. In this study, two scenario areas were evaluated in Tokyo, Japan: Shin-nakano station and Shibuya station surroundings. A 3D model of the Shin-nakano station area is provided at Level of Details 1 (LOD1), where buildings are represented as solids, and all surfaces covering the solid are described as polygon elements. In contrast, the 3D model of the Shibuya station area is provided at LOD2, which allows for more detailed structures than LOD1, enabling the representation of walls, roofs, and floors with different materials. For the LOD1 Shin-nakano model, all materials, including roads, walls, and roofs, were designated as concrete, and the scene file was saved in XML format. In contrast, in the LOD2 Shibuya model, glass is assigned to building walls, concrete to road surfaces, and metal to the overpass surfaces. These materials are determined using the values of the real part of the relative permittivity and conductivity, given by the equations , , where is the carrier frequency, as described in [45]. The specific parameter values for each material and the values of and at the frequency used in this study are listed in Table 2.
@@ 2.2.1 Preparation of 3D Building Data
Metal | 1 | 0 | 0 | 1 |
2.2.2 Generation of Mobility Scenarios
@@
The generation of mobility evaluation scenarios was handled by SUMO (Simulation of Urban MObility) [46]. SUMO is a vehicular traffic simulator capable of realistically replicating vehicle movement by incorporating traffic lights, multiple lanes, speed limits for each road, gradual acceleration and deceleration, and interaction with other vehicles. Based on OpenStreetMap road data, SUMO enables the creation of vehicle mobility scenarios by specifying the driving lanes and their sequence and generating time-series updates of vehicle position coordinates . Although not used in this study, the mobility scenario can also include the velocity vector of the vehicle at each snapshot, which can be utilized to compute the Doppler shift as a key factor affecting fast fading. In both the Shin-nakano and Shibuya scenarios, vehicles temporarily stopped at traffic lights at intersections and made left turns. Additionally, in the narrow residential streets of Shin-nakano, where intersections and T-junctions lack traffic lights, vehicles either slow down or make a temporary stop before proceeding straight or making a left turn.
2.2.3 Channel Impulse Response Calculation Using Ray Tracing
@@
Using any ray-tracing simulator, a deterministic and site-specific channel model for mobility can be computed by integrating the mobility scenario generated by SUMO with 3D urban geometry data from OSM or PLATEAU. At arbitrary time intervals, the CIR can be calculated by considering reflections, diffractions, and diffuse scatterings between the gNB, which is fixed at a specific location, and the UE, which moves according to the mobility scenario.
@@
NVIDIA Sionna RT (v0.18.0) is an OSS that leverages GPU acceleration based on TensorFlow to perform fast ray-tracing simulations of propagation environments [28]. Sionna RT supports two channel models defined in 3GPP TR 38.901: the Clustered Delay Line (CDL) model and the Tapped Delay Line (TDL) model [47]. Both express multipath propagation as a “tapped delay line,” but they differ in that the CDL treats each tap as a cluster and retains angular information, whereas the TDL considers only the delay and power of the taps in the time domain. We used isotropic antennas at both gNB and UE and did not evaluate MIMO/beamforming in this study; accordingly, we dispensed with ray/cluster angle information and adopted the simpler TDL model. Sionna RT was used to efficiently conduct ray tracing over hundreds of snapshots throughout the scenario. However, other ray-tracing simulators that can output CIR, such as Wireless InSite and Opal, can also be applied to the same scenario [48, 49, 50, 51, 52]. A performance comparison across specific versions of several major ray-tracing simulators, including Sionna RT, was reported in an earlier study [53]. The resolution of the CIR output from the simulation depends on the 3D model and mobility scenario. In particular, in complex urban environments, achieving a high-fidelity digital twin is feasible only with high-quality spatiotemporal data.
The simulation parameters for ray tracing are listed in Table 3. The common parameters for both radio propagation and system models include the frequency band, center frequency, number of transceivers, number of antennas per transceiver, and sampling rate. In this study, we adopt n77 band (3.3 GHz–4.2 GHz), with a center frequency of 4.01916 GHz. The number of transceivers and antennas per transceiver were set to one. The sampling rate was set to Msps, as supported by the USRP B200 used on the gNB. This sampling rate is also provided as an input argument to the Sionna RT functions when a bandwidth limitation is applied to the delay profile.
@@
For the ray-tracing simulation, an isotropic antenna pattern was assumed, with , that is, 0 dBi for all directions. Reflection is enabled by setting True, and the maximum number of reflections is set to five by defining . When diffraction is set to True, the model considers up to one diffraction per path without a reflection. In other words, the combination of reflections and diffractions was not calculated.
@@
where is the delay index corresponding to the sampling interval . Because the ideal sinc filter violates causality and is non-time-limited, it is necessary to consider from to . However, in practice, these lower and upper limits need to be set to reasonable finite values, which can be computed using the Sionna RT time_lag_discrete_time_channel() function from a given sampling rate and a sufficiently reasonable maximum delay spread µs to be and . Here, for simplicity, we set the lower limit of the considered delay index to .
2.2.4 Convolution of Channel Impulse Response
@@
In the OAI software, wireless emulation can be executed by convolving the precomputed CIR with the baseband signal in real time according to the time progression of the mobility scenario, thereby reflecting the propagation environment in the transmitted and received signals. This process allows the time evolution of the channel owing to changes in the position of the moving object to be reflected in the KPIs obtainable from the OAI.
@@
In the OAI source code, baseband IQ samples corresponding to the lowest part of the PHY layer are handled within the tx_rf() and rx_rf() functions in the nr-ru.c file. The 5G NR frame format at the baseband IQ sample level and CIR convolution method are shown in Fig. 3. In the OAI RAN, the time-domain signal is stored in every slot as txData and rxData, where the baseband signal of CP-OFDM is represented as complex IQ sample data. To incorporate the delay tap effects of a CIR snapshot into the baseband signal, a matrix is generated by shifting the baseband signal sample-by-sample for the total number of CIR taps . Applying the complex CIR vector from the left and computing the inner product results in the received signal with multipath-delay components. However, in our environment, even with the use of the high-performance computation library ‘BLIS’ (described later), real-time convolution of the -tap delay components obtained from ray tracing was not feasible.
@@
Therefore, considering the most significant power first, the CIR is sorted in descending order as , and the rows of the baseband delay matrix are reordered accordingly to form . Here, the number of leading taps to be read was parameterized as to determine the upper limit of the real-time executable. This approach allows the computation to reflect the most significant multipath components first. The convolution operation was implemented within the rx_rf() and tx_rf() functions on the OAI gNB using cblas_cgemm from the high-performance matrix computation library ‘BLIS’ [54]. Specifically, convolution was applied to the baseband signal before cable transmission from the gNB for the downlink (DL) and after cable reception from the UE for the uplink (UL). The function cblas_cgemm multiplies two complex matrices and multiplies the resulting matrix by a complex constant . Matrix is multiplied by another complex constant . It stores the sum of these two products in matrix as follows: . We used this function to perform the following calculations.
@@
where denotes the index of the sampling time.
@@
To reproduce the effects of the CIR on the transmitted and received signals more accurately, it is necessary to perform convolution with a time-varying CIR for each IQ sample, considering the Doppler effect. However, from a computational cost perspective, real-time execution is infeasible without hardware acceleration, such as an FPGA or GPU [56]. In this study, convolution with a time-invariant CIR was performed every 100 ms to achieve real-time E2E wireless emulation on a commercial CPU-equipped server and commercial SDR without using such accelerators. As mentioned earlier, signal processing in OAI operates on a slot basis, meaning that with the current parameter settings, the CIR convolution is performed once every 0.5 ms. That is, the CIR data were updated every 200 iterations.
2.2.5 Wireless Communication Emulation Using OAI
The system parameters configured in the OAI are summarized in the lower part of Table 3. The bandwidth was set to 40 MHz, which is the maximum achievable bandwidth within the constraints of the sampling rate. The subcarrier spacing is set to 30 kHz, corresponding to numerology . The 1,536‑point FFT employed here uses a three‑quarter sampling frequency of the conventional -sample base clock, thereby avoiding unnecessary bandwidth consumption; this energy‑efficient FFT option is described in 3GPP TR 25.814 [57]. OAI supports this FFT length when operating a 40 MHz channel on a USRP B200 configured for a 46.08 Msps sampling rate. Because only one TX and one RX antenna are used, the system operates in a single-input single-output (SISO) configuration. The duplex mode is TDD, and the slot pattern follows ‘DDDSU.’ The special slot format follows a typical pattern with 6 DL symbols, 4 guard symbols, and 4 UL symbols.
3 RESULTS AND DISCUSSION
3.1 RAY-TRACING SIMULATION RESULTS

### v4

@@ II METHODS
This section outlines the architecture and operational workflow of the proposed platform. Subsection II-A describes the modular architecture of the OWDT platform, detailing how OSSs and open data are integrated to create a realistic and flexible 5G emulation environment. Subsection II-B explains the overall workflow of the OWDT platform, illustrating how data flow through the system from scenario setup and simulation to near-RT monitoring of KPIs.
II-A SYSTEM ARCHITECTURE
@@
We realized the OWDT by integrating the OAI, which reproduces the 5G NR protocol stack from the PHY layer and above, with NVIDIA Sionna RT v0.18.0, which enables the construction of a deterministic channel model through ray tracing. In this environment, CIRs pre-generated using Sionna RT were applied as propagation channels between the next-generation NodeB (gNB) and UE in the OAI. This allows us to simulate a scenario in which the UE moves within a 3D map while performing 5G communication in a laboratory setup. Furthermore, we can observe how KPIs, such as reference signal received powers (RSRPs), modulation and coding schemes (MCSs), block error rates (BLERs), and throughputs, dynamically change in the near-RT during the scenarios. These measurements were performed using the O-RAN standard KPM v3.00 [40], which is one of the service models supported by FlexRIC [20], a near-RT RIC that provides an xApp SDK framework developed within OAI.
@@
The fundamental architecture of the OWDT is illustrated on the right side of Fig. 1. The hardware components that constitute the architecture are as follows: (i) A COTS server to implement the 5GCN, gNB, and near-RT RIC, along with an SDR device responsible for RF signal processing. The server was equipped with an AMD Ryzen 9 7900X 12-core processor (×24), 32 GiB RAM, and ran Ubuntu 22.04.4 LTS with Linux kernel version 6.8.1-060801-generic. An NI USRP B200 was used as the SDR device and connected via USB. In practice, this setup was realized using Allbesmart’s OAIBOX™ 40 [41]. (ii) A computation server used for ray-tracing-based radio propagation simulation with Sionna RT. The server was equipped with an Intel(R) Xeon(R) Silver 4310 @ 2.10 GHz CPU, 256 GiB RAM, an NVIDIA A100 GPU (80 GB), and ran Ubuntu 22.04.4 LTS. (iii) A COTS laptop and SDR device pair for implementing an NR UE, or alternatively, any UE module that can be physically connected to the gNB via an RF cable. In this study, the Quectel RM500Q-GL UE module was connected to a laptop with an 11th Gen Intel Core i7-11370H @ 3.30 GHz (×8), 15.4 GiB RAM, and Ubuntu 20.04 LTS via USB.
II-B WORKFLOW
@@
II-B1 Preparation of 3D Building Data
To achieve a higher resolution in radio propagation simulations, it is essential to prepare highly detailed 3D building data in terms of size and location and to perform site-specific deterministic evaluations. The 3D urban data provided by PLATEAU, an open data project promoted by the Ministry of Land, Infrastructure, Transport, and Tourism (MLIT) in Japan [42], adopts building height measurements obtained via aerial surveying and generally provides higher accuracy in the height dimension than OpenStreetMap (OSM) [43]. These data are available in the CityGML format, which cannot be directly handled by Sionna RT. Therefore, the open-source 3D computer graphics (CG) software Blender 4.2 [44] was used to convert the data into the PLY format. If OSM is used as 3D building data, it can be imported into Blender using Blosm [45]. The Mitsuba Blender Add-on [46] was used to assign materials to the polygon surfaces of the imported 3D model. In this study, two scenario areas were evaluated in Tokyo, Japan: Shin-nakano station and Shibuya station surroundings. A 3D model of the Shin-nakano station area is provided at Level of Details 1 (LOD1), where buildings are represented as solids, and all surfaces covering the solid are described as polygon elements. In contrast, the 3D model of the Shibuya station area is provided at LOD2, which allows for more detailed structures than LOD1, enabling the representation of walls, roofs, and floors with different materials. For the LOD1 Shin-nakano model, all materials, including roads, walls, and roofs, were designated as concrete, and the scene file was saved in XML format. In contrast, in the LOD2 Shibuya model, glass is assigned to building walls, concrete to road surfaces, and metal to overpass surfaces. These materials are determined using the values of the real part of the relative permittivity and conductivity, given by the equations , , where is the carrier frequency, as described in [47]. The specific parameter values for each material and the values of and at the frequency used in this study are listed in Table III.
@@ II-B1 Preparation of 3D Building Data
| Metal | 1 | 0 | 0 | 1 | ||
II-B2 Generation of Mobility Scenarios
@@
The generation of mobility evaluation scenarios was handled by SUMO (Simulation of Urban MObility) [48]. SUMO is a vehicular traffic simulator capable of realistically replicating vehicle movement by incorporating traffic lights, multiple lanes, speed limits for each road, gradual acceleration and deceleration, and interactions with other vehicles. Based on OSM road data, SUMO enables the creation of vehicle mobility scenarios by specifying the driving lanes and their sequence and generating time-series updates of vehicle position coordinates . Although not used in this study, the mobility scenario can also include the velocity vector of the vehicle at each snapshot, which can be utilized to compute the Doppler shift as a key factor affecting fast fading. In both the Shin-nakano and Shibuya scenarios, vehicles temporarily stopped at traffic lights at intersections and made left turns. Additionally, in the narrow residential streets of Shin-nakano, where intersections and T-junctions lack traffic lights, vehicles either slow down or make a temporary stop before proceeding straight or making a left turn.
II-B3 Channel Impulse Response Calculation Using Ray Tracing
@@
Using any ray-tracing simulator, a deterministic and site-specific channel model for mobility can be computed by integrating the mobility scenario generated by SUMO with 3D urban geometry data from the OSM or PLATEAU. At arbitrary time intervals, the CIR can be calculated by considering reflections, diffractions, and diffuse scatterings between the gNB, which is fixed at a specific location, and the UE, which moves according to the mobility scenario.
@@
NVIDIA Sionna RT (v0.18.0) is an OSS that leverages GPU acceleration based on TensorFlow to perform fast ray-tracing simulations of propagation environments [28]. Sionna RT supports two channel models defined in 3GPP TR 38.901: the Clustered Delay Line (CDL) model and the Tapped Delay Line (TDL) model [49]. Both express multipath propagation as a “tapped delay line,” but they differ in that the CDL treats each tap as a cluster and retains angular information, whereas the TDL considers only the delay and power of the taps in the time domain. We used a single isotropic antenna at both the gNB and the UE and did not evaluate Multiple-Input Multiple-Output (MIMO) or beamforming. Consequently, we did not use ray/cluster angles of arrival or antenna/beam directivity, and instead adopted a simpler TDL model, while also avoiding the increased processing load associated with multi-stream transmission. Sionna RT was used to efficiently conduct ray tracing over hundreds of snapshots throughout the scenario. However, other ray-tracing simulators that can output CIR, such as Wireless InSite and Opal, can also be applied to the same scenario [50, 51, 52, 53, 54]. A performance comparison across specific versions of several major ray-tracing simulators, including Sionna RT, was reported in a previous study [55]. The resolution of the CIR output from the simulation depends on the 3D model and mobility scenarios. In particular, in complex urban environments, achieving a high-fidelity digital twin is feasible only with high-quality spatiotemporal data sets.
@@
The simulation parameters for ray tracing are listed in Table IV. The common parameters for both radio propagation and system models include the frequency band, center frequency, number of transceivers, number of antennas per transceiver, and sampling rate. In this study, we adopt n77 band (3.3 GHz–4.2 GHz), with a center frequency of 4.01916 GHz. The number of transceivers and antennas per transceiver were set to one. The sampling rate was set to Msps, as supported by the USRP B200 used on the gNB. This sampling rate is also provided as an input argument to the Sionna RT functions when a bandwidth limitation is applied to the delay profile.
For the ray-tracing simulation, an isotropic antenna pattern was assumed, with , that is, 0 dBi for all directions. Reflection is enabled by setting True, and the maximum number of reflections is set to five by defining . When diffraction is set to True, the model considers up to one diffraction per path without any reflection. In other words, the combination of reflections and diffractions was not calculated.
@@
where is the delay index corresponding to the sampling interval . Because the ideal sinc filter violates causality and is non-time-limited, it is necessary to consider from to . However, in practice, these lower and upper limits need to be set to reasonable finite values, which can be computed using the Sionna RT time_lag_discrete_time_channel() function from a given sampling rate and a sufficiently reasonable maximum delay spread µs to be and . For simplicity, we set the lower limit of the considered delay index to .
II-B4 Convolution of Channel Impulse Response
@@
In the OAI software, wireless emulation can be executed by convolving the pre-computed CIR with the baseband signal in real time according to the time progression of the mobility scenario, thereby reflecting the propagation environment in the transmitted and received signals. This process allows the time evolution of the channel, owing to changes in the position of the moving object, to be reflected in the KPIs obtainable from the OAI.
@@
In the OAI source code, baseband IQ samples corresponding to the lowest part of the PHY layer are handled within the tx_rf() and rx_rf() functions in the nr-ru.c file. The 5G NR frame format at the baseband IQ sample level and the CIR convolution method are shown in Fig. 3. In the OAI RAN, the time-domain signal is stored in every slot as txData and rxData, where the baseband signal of Cyclic Prefix Orthogonal Frequency Division Multiplexing (CP-OFDM) is represented as complex IQ sample data. To incorporate the delay tap effects of a CIR snapshot into the baseband signal, a matrix is generated by shifting the baseband signal sample-by-sample for the total number of CIR taps . Applying the complex CIR vector from the left and computing the inner product results in the received signal with multipath delay components. However, in our environment, even with the use of the high-performance computation library ‘BLIS’ (described later), real-time convolution of the -tap delay components obtained from ray tracing was not feasible.
@@
Therefore, considering the most significant power first, the CIR is sorted in descending order as , and the rows of the baseband delay matrix are reordered accordingly to form . Here, the number of leading taps to be read is parameterized as to determine the upper limit of the real-time executable. This approach allows the computation to reflect the most significant multipath components first. The convolution operation was implemented within the rx_rf() and tx_rf() functions on the OAI gNB using cblas_cgemm from the high-performance matrix computation library ‘BLIS’ [56]. Specifically, convolution was applied to the baseband signal before cable transmission from the gNB for the downlink (DL) and after cable reception from the UE for the uplink (UL). The function cblas_cgemm multiplies two complex matrices and multiplies the resulting matrix by a complex constant . Matrix is multiplied by another complex constant . It stores the sum of these two products in matrix as follows: . We used this function to perform the following calculations.
@@
where denotes the sampling time index.
@@
To reproduce the effects of the CIR on the transmitted and received signals more accurately, it is necessary to perform convolution with a time-varying CIR for each IQ sample, considering the Doppler effect. However, from a computational cost perspective, real-time execution is infeasible without hardware acceleration, such as FPGA or GPU [58]. In this study, we update the CIR every 100 ms and achieve real-time, E2E wireless emulation using only a commercial CPU-based server and a commercial SDR without any accelerators. As noted earlier, because OAI’s signal processing operates on a slot basis, the convolution with the CIR is executed once every 0.5 ms with the current parameter settings. In other words, the same CIR data is reused for 200 consecutive slots.
II-B5 Wireless Communication Emulation Using OAI
The system parameters configured in the OAI are summarized in the lower part of Table IV. The bandwidth was set to 40 MHz, which is the maximum achievable bandwidth within the constraints of the sampling rate. The subcarrier spacing is set to 30 kHz, corresponding to numerology . The 1,536‑point FFT employed here uses a three-quarter sampling frequency of the conventional -sample base clock, thereby avoiding unnecessary bandwidth consumption; this energy-efficient FFT option is described in 3GPP TR 25.814 [59]. OAI supports this FFT length when operating a 40 MHz channel on a USRP B200 configured for a 46.08 Msps sampling rate. Since the focus of this paper is first to demonstrate that the wireless communication system and the propagation phenomena can be simulated/emulated in a synchronized manner, we simplified the problem formulation as much as possible; the system operates in a single-input single-output (SISO) configuration. The duplex mode is Time Division Duplex (TDD), and the slot pattern follows ‘DDDSU.’ The special slot format follows a typical pattern with 6 DL symbols, 4 guard symbols, and 4 UL symbols.
III RESULTS AND DISCUSSION
III-A RAY-TRACING SIMULATION RESULTS